### v2

# Nonholonomic Robot Parking by Feedback— Part II: Nonmodular, Inverse Optimal, Adaptive, Prescribed/Fixed-Time and Safe Designs

Abstract
For the unicycle system, we provide constructive methods for the design of feedback laws that have one or more of the following properties: being nonmodular and globally exponentially stabilizing, inverse optimal, robust to arbitrary decrease or increase of input coefficients, adaptive, prescribed/fixed-time stabilizing, and safe (ensuring the satisfaction of state constraints). Our nonmodular backstepping feedbacks are implementable with either unidirectional or bidirectional velocity actuation. Thanks to constructing families of strict CLFs for the unicycle, we introduce a general design framework and families of feedback laws for the unicycle, which are inverse optimal with respect to meaningful costs. These inverse optimal feedback laws are endowed with robustness to actuator uncertainty and arbitrarily low input saturation due to the unicycle’s driftlessness. Besides ensuring robustness to unknown input coefficients, we also develop adaptive laws for these unknown coefficients, enabling the achievement of good transient performance with lower initial control effort. Additionally, we develop controllers that achieve stabilization within a user-specified time horizon using two systematic methods: time-dilated prescribed-time design with smooth-in-time convergence to zero of both the states and the inputs and homogeneity-based fixed-time control that provides an explicit bound on the settling time. Finally, with a nonovershooting design we guarantee strictly forward motion without curb violation. This article, along with its Part I, lays a broad constructive design foundation for stabilization of the nonholonomic unicycle.
1 Introduction
For the unicycle system, known from the classic results by Brockett, Ryan, Coron, and Rosier [3, 32, 4] to not be stabilizable by static state feedback in Cartesian coordinates, in a companion paper [39] we introduce a broad framework for designing feedback laws and strict global CLFs for the polar coordinate model. Our framework in [39] is modular, meaning that, by separating the design for the feedback laws for the longitudinal and angular velocity inputs, it allows feedback synthesis using the three principal nonlinear feedback design methods—passivity, backstepping, and integrator forwarding. Building on this foundation, in this article we first develop nonmodular stabilizing controllers inspired by [31] to achieve exponential stability, and then expand the feedback design methodology for the unicycle to advanced stabilization goals: inverse optimality, adaptive stabilization, prescribed-time stabilization, and constraint-compliant (safe) stabilization.
1.1 Related works
Inverse optimality
Classical optimal control suffers from the curse of dimensionality, making both solving and storing solutions of the Hamilton-Jacobi-Bellman (HJB) equation infeasible for high-dimensional systems, motivating Kalman [16] to formulate the inverse optimal control problem for linear systems. Introduced for nonlinear systems in [29], the inverse optimal approach has since been extended in various directions, including stochastic systems [6] and differential games with bounded disturbances [8], later generalized to arbitrary disturbances through input-to-state stabilizing controllers [23]. The first constructive exploration of nonlinear inverse optimal control appears in [35], for feedback laws of the form $(L_{g}V)^{\rm T}=-g^{\rm T}\nabla V$, where $g$ is the input vector field of a system affine in control. This approach has found success in applications such as attitude control of rigid spacecraft [24, 2], as well as in inverse optimal safety [20, 21], yet remains considerably underdeveloped for stabilization of nonholonomic systems. Notably, the work of [7] addresses the stabilization of stochastic nonholonomic systems within the inverse optimal framework; however, establishes inverse optimality only at the subsystem level and without strict CLFs thanks to the inherent properties of stochastic systems.
Unicycle adaptive control with model uncertainties
Adaptive control for nonholonomic systems is largely limited to time-varying, switching, or hybrid methods for stabilization and tracking problems. Results such as [14] achieve global asymptotic stabilization using explicitly time-varying feedback for chained-form systems, while [11] employs logic-based switching to handle parametric uncertainties. Robust and adaptive switching schemes also provide adaptive control laws that overcome Brockett’s obstruction and address model uncertainties [12, 9]. Other works, such as [13], focus on adaptive tracking rather than regulation. The absence of globally strict CLFs historically prevented the construction of time-invariant adaptive feedback laws for unicycles, resulting in a literature dominated by time-varying or switching schemes. The recent development of globally strict CLFs [39, 31] allows for the design of time-invariant adaptive feedback laws.
Prescribed/Fixed-time stabilization
The concepts of fixed-time (FxT) stabilization via nonsmooth Lyapunov inequalities and prescribed-time (PT) stabilization via time-varying feedback are introduced in [30, 34, 36, 19] for a broad class of nonlinear systems, respectively. For unicycle models, FxT and PT stabilization results remain limited; in the context of nonholonomic source seeking, we present a prescribed-time seeking method in [40]. In contrast, finite-time convergence is achieved in Dubins vehicles using deadbeat controllers [25] and in unicycles via sliding-mode methods [38]. In addition, fixed-time algorithms for nonholonomic consensus tracking are developed in [5], and for FxT stabilization of nonholonomic systems in chained form in [33]. This paper develops prescribed-time and fixed-time control laws that guarantee parking of the unicycle within a user-specified finite time.
Safety-critical control
Analytical safety filters enable real-time enforcement of safety constraints without relying on reachability analysis or set approximations. Among these, Control Barrier Function (CBF) methods [41, 1] have become a standard framework for ensuring safety in nonlinear systems, with recent developments in inverse optimal safety filters providing safety with built-in optimality and performance tradeoffs [20, 27]. The simplicity of implementation and rigorous theoretical guarantees enable widespread adoption across robotic platforms, including unicycle models [26, 17]. Preceding the development of CBFs, the nonovershooting control technique [22] offers an alternative analytical route to constraint satisfaction. Unlike optimization-based safety filters, this approach provides feedback laws that inherently enforce strict output constraints without requiring a separate filter layer.
1.2 Contributions and Organization
Building on the modular framework and strict CLFs in the companion paper [39], this work establishes a comprehensive set of nonmodular, inverse optimal, adaptive, prescribed-time, fixed-time, and safety-critical stabilization results for the unicycle system. The main contributions are:
Nonmodular stabilization with strict CLFs
In Section 3 we develop nonmodular stabilization feedback laws, following [31], which achieve global exponential stability with backstepping for both unidirectional and bidirectional velocity actuation.
General inverse optimal control for the unicycle
In Section 4 we exploit our construction of globally strict CLFs in [39] to design an extensive family of inverse optimal controllers, including linear, sublinear, bounded, and “relay-approximating” feedback laws, providing a systematic way to trade off control effort and convergence behavior while preserving global stability. Portions of the results presented without proofs appear in [18].
Gain margin and robustness
We show that the inverse optimal controllers inherit an infinite gain margin, a consequence of the unicycle’s driftless structure. This yields robustness to actuator uncertainty and allows stabilization even under arbitrarily low input saturation levels.
Adaptive control with unknown input coefficients
In Section 5, we design adaptive feedback and update laws that compensate for unknown input coefficients while preserving global asymptotic stability. Our approach leverages the globally strict CLFs in [39], which enable smooth and time-invariant adaptive feedback design. Unlike prior adaptive stabilization results, our design avoids the trajectory zigzagging due to oscillating gains inherent to time-varying methods [14], does not rely on the assumption of known bounds on the uncertain parameters as in [12], and eliminates the potential chattering or excessive oscillations from switching control laws in [11, 9]. Adaptation allows for the initial gains to be low while the state is large, and the gains used when the state becomes small to be larger—enabling good transients at reduced (early) peaks in control effort.
Prescribed/Fixed-time stabilization
In Section 6, by properly scaling the nonmodular stabilizing control laws, we guarantee convergence to the origin in a user defined prescribed/fixed-time, with inputs that are not only bounded but whose values, and derivatives, also go to zero in prescribed/fixed-time. The scaling factor in the developed PT-control laws is time-varying and smooth in time and it is derived based on the temporal transformation approach. In contrast, similar to the use of homogeneous scaling factors in time-varying control design [28], where asymptotic stability is achieved, but only with exponential (infinite-time) convergence, we introduce a homogeneous state-dependent scaling factor that enforces a suitable Lyapunov inequality, thereby guaranteeing fixed-time convergence of the closed-loop system. Consequently, the proposed designs provide the first control laws that guarantee user-prescribed finite-time convergence while exhibiting a less restrictive region of attraction than existing approaches.
Safety design
In Section 7, we add curb-like position constraints on top of the stabilization objective and guarantee curb-avoiding stabilization with strictly forward motion, guaranteeing inherent safety without performance-degrading safety filters.
2 Unicycle in Polar Coordinates
Consider the standard unicycle model given as
$\dot{x}$ $=$ $v\cos(\theta)$ (1a)
$\dot{y}$ $=$ $v\sin(\theta)$ (1b)
$\dot{\theta}$ $=$ $\omega\,$, (1c)
where $(x,y)\in\mathbb{R}^{2}$ is the position of the unicycle in Cartesian coordinates, $\theta\in\mathbb{R}$ is the heading angle, $v$ is the forward velocity input, and $\omega$ is the angular velocity input. We introduce the polar transformation of the system (1) as shown in Table 1, obtaining the unicycle model in polar coordinates:
$\dot{\rho}$ $=$ $-v\cos\gamma$ (2a)
$\dot{\delta}$ $=$ $\frac{v}{\rho}\sin\gamma$ (2b)
$\dot{\gamma}$ $=$ $\frac{v}{\rho}\sin\gamma-\omega\,.$ (2c)
Additionally, we define the following state-spaces for the system:
$\mathcal{S}$ $:=\left\{\rho>0\right\}\times\mathcal{T},$ $\quad\mathcal{T}$ $:=\left\{\delta\in\mathbb{R},\,\gamma\in\mathbb{R}\right\}$ (3)
$\mathcal{S}_{1}$ $:=\left\{\rho>0\right\}\times\mathcal{T}_{1},$ $\quad\mathcal{T}_{1}$ $:=\left\{\delta\in\mathbb{R},\,\left\lvert\gamma\right\rvert<\pi\right\}$ (4)
$\mathcal{S}_{2}$ $:=\left\{\rho>0\right\}\times\mathcal{T}_{2},$ $\quad\mathcal{T}_{2}$ $:=\left\{\left\lvert\delta\right\rvert<\pi,\,\gamma\in\mathbb{R}\right\}$ (5)
$\mathcal{S}_{3}$ $:=\left\{\rho>0\right\}\times\mathcal{T}_{3},$ $\quad\mathcal{T}_{3}$ $:=\left\{\left\lvert\delta\right\rvert<\pi,\,\left\lvert\gamma\right\rvert<\pi\right\}\,.$ (6)
Figure 1: Unicycle orientation $(x,y,\theta)$ relative to the goal state $(0,0,0)$, and the corresponding polar coordinate transformation $(\rho,\gamma,\delta)$.
$\rho=\sqrt{x^{2}+y^{2}}$ Distance to origin
$\delta=\text{{\rm atan2}}(y,x)+\pi$ Polar angle
$\gamma=\text{{\rm atan2}}(y,x)-\theta+\pi$ Line-of-sight angle
Table 1: Polar coordinates and their expressions in terms of Cartesian coordinates. The transformation $(x,y,\theta)\mapsto(\rho,\delta,\gamma)$ is discontinuous on $x<0,y=0$ and not defined at $x=y=0$. If the target is at $(x^{*},y^{*},\theta^{*})\neq 0$, the transformation generalizes to $\rho=\sqrt{(x-x^{*})^{2}+(y-y^{*})^{2}},\delta=\text{{\rm atan2}}(y-y^{*},x-x^{*})-\theta^{*}+\pi,\gamma=\delta-\theta+\theta^{*}$.
3 Nonmodular Unicycle Stabilization
The companion paper [39] develops a modular framework for designing stabilizing controllers and global strict CLFs for the unicycle. Modularity allows multiple design methods (passivity, integrator forwarding, backstepping) but guarantees only global asymptotic stability (GAS), not global exponential stability (GES). Notably, sacrificing modularity enables GES via a backstepping design [31] that jointly determines the forward velocity and steering inputs in the first step for the $(\rho,\delta)$-subsystem. This, however, comes at the cost of higher design complexity and the lack of extension beyond backstepping to passivity or forwarding methods. Despite these limitations, we develop two backstepping extensions of the nonmodular approach in [31].
3.1 Unidirectional GES BAR-FLi
In many practical systems, such as fixed-wing aircraft or guided missiles, reversing the forward velocity is infeasible. Motivated by this, we first present an extension for the GES controllers from [31] to prevent crossing in front of the target under the constraint of unidirectional forward velocity.
Theorem 1 (Unidirectional BAR-FLi).
For the system (2), consider the feedback laws
$v$ $=$ $k_{1}\sigma(\delta)\rho$ (7)
$\omega$ $=$ $\dfrac{v}{\rho}\sin(\gamma)+\tilde{\omega}$ (8)
$\tilde{\omega}$ $=$ $k_{4}z-k_{3}\rho^{2}\sigma(\delta)\frac{\partial\psi(z,\gamma)}{\partial\gamma}+k_{3}\frac{V_{\delta}(\delta)}{\phi(\delta)}\sigma(\delta)\psi(z,\gamma)$ (9)
$+\frac{k_{1}k_{2}}{\sigma^{2}(\delta)}\phi^{\prime}(\delta)\left(\sigma(\delta)\psi(z,\gamma)z-\phi(\delta)\right)\,$,
where
$V_{\delta}(\delta)$ $=$ $4\tan^{2}\frac{\delta}{2}$ (10)
$\phi(\delta)$ $=$ $\sin\delta$ (11)
$\sigma(\delta)$ $=$ $\sqrt{1+k_{2}^{2}\phi^{2}(\delta)}$ (12)
$z$ $=$ $\gamma+\arctan\left(k_{2}\phi(\delta)\right)$ (13)
$\psi(r,s)$ $=$ $\frac{\sin(r-s)+\sin(s)}{r}\,$, (14)
and $k_{1},k_{2},k_{3},k_{4}>0$. This feedback renders the point $\rho=\delta=\gamma=0$ globally exponentially stable (GES) on $\mathcal{S}_{2}$, i.e., there exist $K\geq 1$ and $\lambda>0$ such that, for all $t\geq 0$, it holds that $|(\rho(t),\delta(t),\gamma(t))|_{\mathcal{S}_{2}}\leq K|(\rho_{0},\delta_{0},\gamma_{0})|_{\mathcal{S}_{2}}e^{-\lambda(t-t_{0})}$. Furthermore,
$V=\rho^{2}+V_{\delta}+q^{2}\left[\gamma+\arctan\left(k_{2}\phi(\delta)\right)\right]^{2}\,$, (15)
with $q=\sqrt{k_{1}/k_{3}}$ is a globally strict CLF for (2) on $\mathcal{S}_{2}$ with respect to the input pair $(v/\rho,\omega)$ in the sense of [39, Def. 2].
Proof. First, we note that $\sigma(\delta)\cos(z-\gamma)=1$ and $\sigma(\delta)\sin(z-\gamma)=\phi(\delta)$. With the backstepping transformation (13) and substituting (7) into (2a) and (2b) yields
$\dot{\rho}$ $=$ $-k_{1}\rho-k_{1}\rho\sigma(\delta)\frac{\partial\psi(z,\gamma)}{\partial\gamma}z$ (16)
$\dot{\delta}$ $=$ $-k_{1}k_{2}\phi(\delta)+k_{1}\sigma(\delta)\psi(z,\gamma)z\,.$ (17)
Moreover, with (8), we get
$\dot{z}=\frac{k_{1}k_{2}}{\sigma^{2}(\delta)}\phi^{\prime}(\delta)\left(\sigma(\delta)\psi(z,\gamma)z-k_{2}\phi(\delta)\right)-\bar{\omega}\,.$ (18)
Then, taking the time derivative of (15) gives
$\dot{V}=$ $-2k_{1}\rho^{2}-2k_{1}k_{2}V_{\delta}(\delta)$
$+2q^{2}z\biggl[-k_{3}\rho^{2}\sigma(\delta)\frac{\partial\psi(z,\gamma)}{\partial\gamma}+k_{3}\frac{V_{\delta}(\delta)}{\phi(\delta)}\sigma(\delta)\psi(z,\gamma)$
$+\frac{k_{1}k_{2}}{\sigma^{2}(\delta)}\phi^{\prime}(\delta)\left(\sigma(\delta)\psi(z,\gamma)z-k_{2}\phi(\delta)\right)-\bar{\omega}\biggr]\,$, (19)
which, with (9), yields
$\dot{V}=-2k_{1}\rho^{2}-2k_{1}k_{2}V_{\delta}(\delta)-2k_{4}q^{2}z^{2}\leq-\underline{c}V\,$, (20)
where $\underline{c}=\min\{2k_{1},2k_{1}k_{2},2k_{4}q^{2}\}>0$. Hence, (15) is a strict CLF for (2), and by the comparison principle $|(\rho(t),\delta(t),\gamma(t))|_{\mathcal{S}_{2}}\leq K|(\rho_{0},\delta_{0},\gamma_{0})|_{\mathcal{S}_{2}}e^{-\lambda(t-t_{0})}$, where $K(k_{2})>1$ and $\lambda=\underline{c}/2$. ∎
3.2 Bidirectional GES Backstepping
When the vehicle is allowed to reverse the forward velocity, it would be more efficient to exploit this capability when needed, especially in parking scenarios. The following Theorem exploits this and achieves GES. Fig. 2 compares the trajectory generated by the controller from Theorem 2 with that of [31, Sec. III.A].
Theorem 2 (Bidirectional Backstepping).
For the system (2), consider the feedback laws
$v$ $=$ $k_{1}\rho\sigma(\delta)\cos\gamma$ (21)
$\omega$ $=$ $\dfrac{v}{\rho}\sin\gamma+\tilde{\omega}$ (22)
$\tilde{\omega}$ $=$ $k_{4}z-k_{3}\rho^{2}\sigma(\delta)\psi_{2}(2z,2\gamma)+2k_{3}\delta\sigma(\delta)\psi(2z,2\gamma)$ (23)
$+\frac{k_{1}k_{2}}{\sigma^{2}(\delta)}\left(\sigma(\delta)\psi(2z,2\gamma)z-k_{2}\delta\right)\,$,
where
$\sigma(\delta)$ $=$ $\sqrt{1+(2k_{2}\delta)^{2}}$ (24)
$z$ $=$ $\gamma+\frac{1}{2}\arctan\left(2k_{2}\delta\right)$ (25)
$\psi_{2}(r,s)$ $=$ $\frac{\partial\psi(r,s)}{\partial s}\,$, (26)
with $\psi(r,s)$ defined in (14) and $k_{1},k_{2},k_{3},k_{4}>0$. This feedback renders the point $\rho=\delta=\gamma=0$ globally exponentially stable (GES) on $\mathcal{S}$, i.e., there exist $K\geq 1$ and $\lambda>0$ such that, for all $t\geq 0$, it holds that $|(\rho(t),\delta(t),\gamma(t))|_{\mathcal{S}}\leq K|(\rho_{0},\delta_{0},\gamma_{0})|_{\mathcal{S}}e^{-\lambda(t-t_{0})}$. Furthermore,
$V=\rho^{2}+\delta^{2}+q^{2}\left[\gamma+\frac{1}{2}\arctan\left(2k_{2}\delta\right)\right]^{2}\,$, (27)
with $q=\sqrt{k_{1}/k_{3}}$ is a globally strict CLF for (2) on $\mathcal{S}$ with respect to the input pair $(v/\rho,\omega)$ in the sense of [39, Def. 2].
Proof. With the backstepping transformation (25) and substituting (21) into (2a) and (2b) yields
$\dot{\rho}$ $=$ $-\frac{1}{2}k_{1}\rho(1+\sigma(\delta))-k_{1}\rho\sigma(\delta)\psi_{2}(2z,2\gamma)z$ (28)
$\dot{\delta}$ $=$ $-k_{1}k_{2}\delta+k_{1}\sigma(\delta)\psi(2z,2\gamma)z\,.$ (29)
Moreover, with (22), we get
$\dot{z}=\frac{k_{1}k_{2}}{\sigma^{2}(\delta)}(\sigma(\delta)\psi(2z,2\gamma)z-k_{2}\delta)-\bar{\omega}\,.$ (30)
Then, taking the time derivative of (27) gives
$\dot{V}=$ $-k_{1}\rho^{2}\left(1+\sigma(\delta)\right)-2k_{1}k_{2}\delta^{2}$
$+2q^{2}z\biggl[-k_{3}\rho^{2}\sigma(\delta)\psi_{2}(2z,2\gamma)+k_{3}\sigma(\delta)\psi(2z,2\gamma)$
$+\frac{k_{1}k_{2}}{\sigma^{2}(\delta)}\left(\sigma(\delta)\psi(2z,2\gamma)z-k_{2}\delta\right)-\bar{\omega}\biggr]\,$, (31)
which, with (23), yields
$\dot{V}=-k_{1}\rho^{2}-2k_{1}k_{2}\delta^{2}-2k_{4}q^{2}z^{2}\leq-\underline{c}V\,$, (32)
where $\underline{c}=\min\{k_{1},2k_{1}k_{2},2k_{4}q^{2}\}>0$. Hence, (27) is a strict CLF for (2), and by the comparison principle $|(\rho(t),\delta(t),\gamma(t))|_{\mathcal{S}}\leq K|(\rho_{0},\delta_{0},\gamma_{0})|_{\mathcal{S}}e^{-\lambda(t-t_{0})}$, where $K(k_{2})>1$} and $\lambda=\underline{c}/2$}. ∎
Figure 2: Comparison of the control law from Theorem 2 (blue) and [31, Sec. III.A] (red) with initial conditions $(\rho_{0},\delta_{0},\gamma_{0})=(1,-4\pi/5,\pi)$, showing that restricting to unidirectional forward velocity yields a less efficient trajectory to the target.
4 Inverse Optimal Design
We now turn to the design of optimal controllers. The inescapable ‘curse of dimensionality’ of classical optimal control is to an equal degree about storing the results as it is about solving the HJB PDE. For instance, on devices with less than a terabyte of storage, there is not enough space to store the HJB solution on a uniformly quantized grid for a robotic manipulator with more than 3DOF. Instead, in this section, we derive a family of inverse optimal controllers for the unicycle based on the numerous CLFs available. It is important to note that the derived controllers fall into the category of damping $L_{g}V$ controllers (i.e., see [15], [37, Prop. 5.9.1]) that immediately satisfy the Jurdjevic-Quinn conditions and globally stabilize the unicycle.
4.1 Alternative controllers: $L_{g}V$ (gradient) designs
We return to the unicycle model (2). If we take the $L_{g}V$ approach in its vanilla form, we run into a problem: $g_{1}(\rho,\gamma)$ has a singularity at $\rho=0$. We circumvent this singularity in a way that is not just mathematically resourceful but physically meaningful. We introduce a change of velocity control variable from $v$ to $v/\rho$, where $v/\rho$ is a new control input to be designed, and where the velocity feedback $v$ is forced to vanish at the target location, $\rho=0$. It is to be expected that the velocity be proportional to the distance from the target—the vehicle needs to slow down in order to stop at the target.
Having replaced the input pair $(v,\omega)$ by $(v/\rho,\omega)$, we now have a new input vector field $\bar{g}=[\rho g_{1},g_{2}]$, with which $L_{\bar{g}}V$ is “desingularized.” From (2), the vector fields are given by
$\quad g_{1}$ $\coloneqq\dfrac{1}{\rho}\bar{g}_{1}\coloneqq\dfrac{1}{\rho}{\begin{bmatrix}-\rho\cos\gamma\\ \sin\gamma\\ \sin\gamma\end{bmatrix}}\,,\ \ g_{2}:={\begin{bmatrix}0\\ 0\\ -1\end{bmatrix}}\,$. (33)
And we are now ready to study the design of controllers that employ $L_{\bar{g}}V$, as well as their stabilizing and inverse optimality properties.
As we shall see shortly, the inverse optimal design will rely on the property that
$L_{\bar{g}}V(\Xi)=0$ only at $\Xi=0$, where $\Xi:=(\rho,\gamma,\delta)$,
which in turn implies that
$|L_{\bar{g}}V(\Xi)|$ is positive definite in $\Xi$.
This property holds for the following reason. For systems affine in control, $\dot{x}=f(x)+g(x)u$, a positive definite radially unbounded differentiable function $V(x)$ is a CLF if and only if the following implication holds: for all $x\neq 0$ where $L_{g}V(x)=0$ it holds that $L_{f}V(x)<0$. When $f(x)\equiv 0$, this implication gives that $L_{g}V(x)\neq 0$ for all $x\neq 0$, namely, that $L_{g}V(x)=0$ only at $x=0$. For instance, we establish that all Lyapunov functions in [39, Thm. 1] are CLFs in the sense of [39, Def. 2] with the Genova feedback law.
While our theorems establish that we construct Lyapunov functions $V(\rho,\delta,\gamma)$ such that $|L_{\bar{g}}V(\Xi)|$ is positive definite in $\Xi$, it is not radially unbounded. Radial unboundedness can be achieved with a modification of the CLF, which is described in the final segment of the proof of [23, Thm. 3.2].
As an illustration of what the Lie derivatives $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$ look like, we compute them for the Genova CLF in [39, Eq. (25) with (13) and (24)], with unity gains:
$-L_{\bar{g}_{1}}V/2$ $=\rho^{2}\cos\gamma-\sin\gamma\left(\delta^{2}+\gamma^{2}+4\right)(\delta+\gamma)$ (34a)
$-L_{g_{2}}V/2$ $=\left(\delta^{2}+\gamma^{2}+2\right)\gamma+(\delta+\gamma)\,.$ (34b)
Similarly, the Lie derivatives $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$ for the GloBa CLF in [39, Eq. (37) with (13) and (38)], with unity gains, yield
$-L_{\bar{g}_{1}}V/2$ $=\rho^{2}\cos\gamma-\sin\gamma\left[\delta+\left(1+\frac{1}{1+4\delta^{2}}\right)z\right]$ (35a)
$-L_{g_{2}}V/2$ $=z\,.$ (35b)
where $z=\gamma+\frac{1}{2}\arctan(2\delta)$. In both cases, the introduction of the change of control variable from $v$ to $v/\rho$ has successfully “desingularized” the Lie derivative expressions as seen in (34a) and (35a).
4.2 Basic quadratic inverse optimality
Before presenting the general inverse optimality methodology in Section 4.3, we give a particular example of what we are after. For any of our previously designed CLFs $V(\rho,\delta,\gamma)$, all controllers of the form
$v^{*}$ $=-\rho\varepsilon_{1}^{2}L_{\bar{g}_{1}}V$
$:=\varepsilon_{1}^{2}\rho\left[\dfrac{\partial V}{\partial\rho}\rho\cos\gamma-\left(\dfrac{\partial V}{\partial\delta}+\dfrac{\partial V}{\partial\gamma}\right)\sin\gamma\right]$ (36a)
$\omega^{*}$ $=-\varepsilon_{2}^{2}L_{g_{2}}V:=\varepsilon_{2}^{2}\dfrac{\partial V}{\partial\gamma}\,$, (36b)
for all $\varepsilon_{1},\varepsilon_{2}>0$, are not only globally asymptotically stabilizing but are the minimizers of the parametrized costs
$J=\int_{0}^{\infty}\Biggl[l(\rho,\delta,\gamma)+\left(\dfrac{v}{\varepsilon_{1}\rho}\right)^{2}+\left(\dfrac{\omega}{\varepsilon_{2}}\right)^{2}\Biggr]dt\,$, (37)
where
$l(\rho,\delta,\gamma)=\left(\varepsilon_{1}L_{\bar{g}_{1}}V\right)^{2}+\left(\varepsilon_{2}L_{g_{2}}V\right)^{2}\,.$ (38)
Such costs are meaningful in the sense of imposing (i.) a zero penalty when all the state variables $\rho,\delta,\gamma$ and both of the ($\rho$-weighted) control inputs $v/\rho$ and $\omega$ are zero, and (ii.) a positive penalty when any of the states and either of the (weighted) controls are nonzero. Note that when $\rho=0$, the forward velocity input is subjected to an infinite penalty, which is desirable, as it discourages movement away from the positional origin once it is reached.
It is relevant to examine the complexity of the $L_{g}V$ controllers as compared to their counterparts used to derive the CLFs. To that end, we focus on the GloBa controller and note that the inverse optimal steering feedback (36b) with (35b) is much simpler than the non-optimal GloBa feedback [39, Eq. (7) with (41)]. But, in contrast, compared to the basic forward velocity feedback $v=\rho\cos\gamma$, the inverse optimal forward velocity feedback (36) with (35a), given by $v^{*}=\rho^{3}\cos\gamma-\rho\sin\gamma\left[\delta+\left(1+\frac{1}{1+4\delta^{2}}\right)z\right]$ is considerably more complex and depends not only on $(\rho,\gamma)$ but also on $\delta$ (when the vehicle points sidewise relative to the target, the velocity is nonzero).
4.3 General inverse optimal designs
To generalize the inverse optimality result from the end of Section 4.2, we recall the Legendre–Fenchel transform, defined as follows.
Definition 1.
(Legendre-Fenchel transform). Let $\eta$ be a class $\mathcal{K}_{\infty}[0,a)$ function whose derivative $\eta^{\prime}$ is also a class $\mathcal{K}_{\infty}[0,a)$ function where $a>0$ is finite or infinite. The mapping
$\ell\eta(r)=\int_{0}^{r}(\eta^{\prime})^{-1}(s)ds$ (39)
represents the Legendre-Fenchel transform, where $(\eta^{\prime})^{-1}(r)$ stands for the inverse function of $d\eta(r)/dr$.
The transformation has useful properties, given in [23, Lemma A1]. Also, to minimize notational repetition, we use the subscript $i$ in place of the subscripts $1$ and $2$, corresponding to the vector fields $\bar{g}_{1}$ and $g_{2}$ and their associated control inputs $v/\rho$ and $\omega$.
Theorem 3.
Consider the system (2) rewritten as
$\dot{\Xi}=\bar{g}_{1}(\Xi)\frac{v}{\rho}+g_{2}(\Xi)\omega$ (40)
where $\Xi\coloneqq[\rho,\delta,\gamma]^{\top}$ and
$\bar{g}_{1}=\begin{bmatrix}-\rho\cos\gamma\\ \sin\gamma\\ \sin\gamma\end{bmatrix},\quad g_{2}=\begin{bmatrix}0\\ 0\\ -1\end{bmatrix}\,$. (41)
For any $\eta_{i}\in\mathcal{K}_{\infty}[0,a_{i})$ such that also $\eta_{i}^{\prime}\in\mathcal{K}_{\infty}[0,a_{i})$ where $a_{i}>0$ is finite or infinite, and for any continuous positive scalar-valued functions $\varepsilon_{1}(\rho,\delta,\gamma)$ and $\varepsilon_{2}(\rho,\delta,\gamma)$, the cost functional
$J=\int_{0}^{\infty}\Biggl[l(\rho,\delta,\gamma)+\eta_{1}\left(\frac{|v|}{\varepsilon_{1}\rho}\right)+\eta_{2}\left(\frac{|\omega|}{\varepsilon_{2}}\right)\Biggr]dt\,$, (42)
where
$l(\rho,\delta,\gamma)=\ell\eta_{1}(\varepsilon_{1}|\nu_{1}|)+\ell\eta_{2}(\varepsilon_{2}|\nu_{2}|)\,$, (43)
is minimized by the feedback law
$v^{*}$ $=-\rho\varepsilon_{1}(\eta_{1}^{\prime})^{-1}(\varepsilon_{1}|\nu_{1}|)\mathrm{sgn}(\nu_{1})$ (44a)
$\omega^{*}$ $=-\varepsilon_{2}(\eta_{2}^{\prime})^{-1}(\varepsilon_{2}|\nu_{2}|)\mathrm{sgn}(\nu_{2})\,$, (44b)
with $\nu_{1}(\Xi)\coloneqq L_{\bar{g}_{1}}V$ and $\nu_{2}(\Xi)\coloneqq L_{g_{2}}V$ denoted for the CLFs given by [39, Thms. 1-8], and for all initial conditions on the respective state spaces $\mathcal{S},\mathcal{S}_{1},\mathcal{S}_{2},$ and $\mathcal{S}_{3}$. Additionally, the feedback law
$v$ $=-\rho\varepsilon_{1}\frac{\ell\eta_{1}(\varepsilon_{1}|\nu_{1}|)}{\varepsilon_{1}|\nu_{1}|}\mathrm{sgn}(\nu_{1})$ (45a)
$\omega$ $=-\varepsilon_{2}\frac{\ell\eta_{2}(\varepsilon_{2}|\nu_{2}|)}{\varepsilon_{2}|\nu_{2}|}\mathrm{sgn}(\nu_{2})\,$, (45b)
is continuous in $(\nu_{1},\nu_{2})$ and renders the point $\rho=\delta=\gamma=0$ of the system (40) GAS on the respective state spaces $\mathcal{S},\mathcal{S}_{1},\mathcal{S}_{2},$ and $\mathcal{S}_{3}$, in accordance to [39, Def. 1].
Proof. Step 1: (45) is continuous in $(\nu_{1},\nu_{2})$ and stabilizes (40). First, from Definition 1 and applying the Leibniz rule yields $(\ell\eta)^{\prime}(r)=(\eta^{\prime})^{-1}(r)$. Then, applying L’Hôpital’s rule, we get
$\lim_{r\rightarrow 0}\frac{\ell\eta(r)}{r}=\lim_{r\rightarrow 0}\frac{(\eta^{\prime})^{-1}(r)}{1}\,.$ (46)
Since $\eta^{\prime}(r)$ is a class $\mathcal{K}_{\infty}$ function, its inverse $(\eta^{\prime})^{-1}(r)$ is also a class $\mathcal{K}_{\infty}$ function, which by definition is zero at $r=0$. Thus, the limit in (46) is zero and the feedback law (45) is continuous in $(\nu_{1},\nu_{2})$.
Secondly, plugging in (45) into $\dot{V}$ yields
[…]
$=-\ell\eta_{1}(\varepsilon_{1}|\nu_{1}|)-\ell\eta_{2}(\varepsilon_{2}|\nu_{2}|)<0\,.$ (47)
By [23, Lemma A1], $\ell\eta_{i}$ are class $\mathcal{K}_{\infty}$ functions, and since $V$ is a strict CLF, it follows by definition that $\nu_{1}(\Xi)=0$ and $\nu_{2}(\Xi)=0$ if and only if $\Xi=0$. Hence, (47) holds, and the point $\rho=\delta=\gamma=0$ is GAS under each CLF on its respective state space when using the feedback law (45).
Step 2: (44) stabilizes (40). Plugging in (44) into $\dot{V}$ yields
[…]
$=-\varepsilon_{1}|\nu_{1}|(\eta_{1}^{\prime})^{-1}(\varepsilon_{1}|\nu_{1}|)-\varepsilon_{2}|\nu_{2}|(\eta_{2}^{\prime})^{-1}(\varepsilon_{2}|\nu_{2}|)\,.$ (48)
From [23, Lemma A1], we see that $r(\eta^{\prime})^{-1}(r)=\ell\eta(r)+\eta((\eta^{\prime})^{-1}(r))$ and (48) can be rewritten as
[…]
Thus, the point $\rho=\delta=\gamma=0$ is GAS under each CLF on its respective state space when using the feedback law (44).
Step 3: (44) minimizes (42). First, we rearrange (42) and consider the fact that $\lim_{t\rightarrow\infty}V(\Xi(t))=0$, yielding
$J$ $=\int_{0}^{\infty}\biggl[\ell\eta_{1}(\varepsilon_{1}|\nu_{1}|)+\ell\eta_{2}(\varepsilon_{2}|\nu_{2}|)$
$\qquad\qquad\qquad+\eta_{1}\left(\frac{|v|}{\varepsilon_{1}\rho}\right)+\eta_{2}\left(\frac{|\omega|}{\varepsilon_{2}}\right)+\dot{V}-\dot{V}\biggr]dt$
$=V(\Xi(0))-\lim_{t\rightarrow\infty}V(\Xi(t))$
$\qquad+\int_{0}^{\infty}\biggl[\ell\eta_{1}(\varepsilon_{1}|\nu_{1}|)+\ell\eta_{2}(\varepsilon_{2}|\nu_{2}|)$
$\qquad\qquad\qquad+\eta_{1}\left(\frac{|v|}{\varepsilon_{1}\rho}\right)+\eta_{2}\left(\frac{|\omega|}{\varepsilon_{2}}\right)+\nu_{1}\frac{v}{\rho}+\nu_{2}\omega\biggr]dt$
$=V(\Xi(0))+\int_{0}^{\infty}\biggl[\ell\eta_{1}(\varepsilon_{1}|\nu_{1}|)+\eta_{1}\left(\frac{|v|}{\varepsilon_{1}\rho}\right)+\nu_{1}\frac{v}{\rho}$
$\qquad\quad\qquad+\ell\eta_{2}(\varepsilon_{2}|\nu_{2}|)+\eta_{2}\left(\frac{|\omega|}{\varepsilon_{2}}\right)+\nu_{2}\omega\biggr]dt\,.$ (50)
Using [10, Thm. 156] and the fact that $\rho>0$ for all $\mathcal{S},\mathcal{S}_{1},\mathcal{S}_{2},$ and $\mathcal{S}_{3}$, we get the following upper bounds
$-\nu_{1}\frac{v}{\rho}$ $=\left(\frac{v}{\varepsilon_{1}\rho}\right)\left(-\varepsilon_{1}\nu_{1}\right)\leq\eta_{1}\left(\frac{|v|}{\varepsilon_{1}\rho}\right)+\ell\eta_{1}(\varepsilon_{1}|\nu_{1}|)$ (51)
$-\nu_{2}\omega$ $=\left(\frac{\omega}{\varepsilon_{2}}\right)\left(-\varepsilon_{2}\nu_{2}\right)\leq\eta_{2}\left(\frac{|\omega|}{\varepsilon_{2}}\right)+\ell\eta_{1}(\varepsilon_{2}|\nu_{2}|)\,.$ (52)
Then, multiplying both sides with 1-1 yields
$\nu_{1}\frac{v}{\rho}$ $\geq-\eta_{1}\left(\frac{|v|}{\varepsilon_{1}\rho}\right)-\ell\eta_{1}(\varepsilon_{1}|\nu_{1}|)$ (53)
$\nu_{2}\omega$ $\geq-\eta_{2}\left(\frac{|\omega|}{\varepsilon_{2}}\right)-\ell\eta_{1}(\varepsilon_{2}|\nu_{2}|)\,$, (54)
and we observe that the cost functional (50) achieves a minimum when (53) and (54) are equalities. The result from [10, Thm. 156] states that (53) and (54) are equalities if and only if
$x_{1}=(\eta_{1}^{\prime})^{-1}(|y_{1}|)\frac{y_{1}}{|y_{1}|},\;\;\text{and}\;\;x_{2}=(\eta_{2}^{\prime})^{-1}(|y_{2}|)\frac{y_{2}}{|y_{2}|},$ (55)
where $x_{1}\coloneqq\frac{v}{\varepsilon_{1}\rho},\;y_{2}\coloneqq-\varepsilon_{1}\nu_{1},\;x_{2}\coloneqq\frac{\omega}{\varepsilon_{2}},$ and $y_{2}\coloneqq-\varepsilon_{2}\nu_{2}$, which implies
$\frac{v}{\varepsilon_{1}\rho}$ $=-(\eta_{1}^{\prime})^{-1}(\varepsilon_{1}|\nu_{1}|)\frac{\varepsilon_{1}\nu_{1}}{\varepsilon_{1}|\nu_{1}|}$
$\implies v=-\rho\varepsilon_{1}(\eta_{1}^{\prime})^{-1}(\varepsilon_{1}|\nu_{1}|)\mathrm{sgn}(\nu_{1})=v^{*}$ (56)
$\frac{\omega}{\varepsilon_{2}}$ $=-(\eta_{2}^{\prime})^{-1}(\varepsilon_{2}|\nu_{2}|)\frac{\varepsilon_{2}\nu_{2}}{\varepsilon_{2}|\nu_{2}|}$
$\implies\omega=-\varepsilon_{2}(\eta_{2}^{\prime})^{-1}(\varepsilon_{2}|\nu_{2}|)\mathrm{sgn}(\nu_{2})=\omega^{*}\,.$ (57)
Thus, the inequalities (53) and (54) are equalities if and only if $v=v^{*}$ and $\omega=\omega^{*}$, which yields
$\nu_{1}\frac{v}{\rho}$ $=-\eta_{1}\left(\frac{|v|}{\varepsilon_{1}\rho}\right)-\ell\eta_{1}(\varepsilon_{1}|\nu_{1}|)$ (58a)
$\nu_{2}\omega$ $=-\eta_{2}\left(\frac{|\omega|}{\varepsilon_{2}}\right)-\ell\eta_{1}(\varepsilon_{2}|\nu_{2}|)\,.$ (58b)
Substituting (58) into (50) yields $J=V(x(0))$. Thus, (44) minimizes (42) for each CLF on its respective state space. ∎
While the $L_{g}V$ controllers achieve optimality, they do not retain the local exponential stabilizing property of the basic stabilizing controllers, showcased in Section X of [39]. This is evident, for example, from the $L_{\bar{g}_{1}}V$ expressions (34a) and (35a), which are locally quadratic in the state near the origin, and therefore cannot be exponentially stabilizing. Such a linearity-lacking dependence is the result of the change of control variable from $v$ to $v/\rho$ and is necessitated by the singularity of the unicycle model at $\rho=0$.
4.4 Specific inverse optimal choices
$\eta(r)$ $(\eta^{\prime})^{-1}(r)$ $\dfrac{\ell\eta(r)}{r}$
$\dfrac{r^{2}}{2}$ $r$ $\dfrac{r}{2}$
$\cosh(r)-1$ $\operatorname{arcsinh}(r)$ $\operatorname{arcsinh}(r)-\dfrac{r}{\sqrt{r^{2}+1}+1}$
$\ln(\cos(r))$ $\arctan(r)$ $\arctan(r)-\dfrac{\ln(1+r^{2})}{2r}$
$\dfrac{e}{e^{{1}/{r}}-e}$ $\dfrac{1}{1+\ln\left(1+1/r\right)}$ $\dfrac{1}{r}\displaystyle\int\dfrac{\text{d}r}{1+\ln\left(1+1/r\right)}$
Table 2: The cost-on-control functions $\eta(\cdot)$ and the resulting expressions that define the inverse optimal and stabilizing $L_{g}V$ controllers.
(a) $\eta(r)=r^{2}/2$ in (black), $(\eta^{\prime})^{-1}(r)=r$ in (red) and $\frac{\ell\eta(r)}{r}=r/2$ in (blue).
(b) $\eta(r)=\cosh(r)-1$ in (black), $(\eta^{\prime})^{-1}(r)={\rm arcsinh}(r)$ in (red) and $\frac{\ell\eta(r)}{r}={\rm arcsinh}(r)-r/(1+\sqrt{r^{2}+1})$ in (blue).
(c) $\eta(r)=-\ln(\cos(r))$ in (black), $(\eta^{\prime})^{-1}(r)=\arctan(r)$ in (red) and $\frac{\ell\eta(r)}{r}=\arctan(r)-\ln(1+r^{2})/2r$ in (blue).
(d) $\eta(r)=e/(e^{1/r}-e)$ in (black), $(\eta^{\prime})^{-1}(r)=1/(1+\ln(1+1/r))$ in (red) and $\frac{\ell\eta(r)}{r}=\frac{1}{r}\int\frac{{\rm d}r}{1+\ln(1+1/r)}$ in (blue).
Figure 3: Cost-on-control functions from Table 2.
The design space of the stabilizer (45) and its optimal counterpart (44) includes many parameters and infinitely many CLF choices. An exhaustive exploration is impossible, but one can explore qualitatively the tradeoff between the control effort and the related cost on the state.
We present four examples based on different choices of the function $\eta$, each leading to a distinct inverse optimal controller. Example 4.4 yields a controller linear in $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$, already introduced in Section 4.2, Example 4.5 produces a sublinear relation of control with $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$, and Examples 4.6 and 4.7 result in bounded control laws. While Theorem 3 does not require uniformity in the choice of $\eta_{1}$ and $\eta_{2}$, the following examples use the same function for both to highlight each behavior, and the subscript on $\eta(r)$ is omitted henceforth. The four choices of $\eta$ are summarized in Table 2, while Figure 3 illustrates how the functions increase with respect to the input, reflecting the corresponding rise in control effort cost.
Additionally, since inverse optimal redesign relies on $L_{g}V$-based controllers, different CLFs naturally lead to different system behaviors. Hence, we conduct simulation comparisons between the examples using a particular GloBa CLF in the following corollary derived from [39, Thm. 7]:
Corollary 1.
For the system (2), the composite Lyapunov function
$V(\rho,\delta,\gamma)=\sqrt{1+k_{1}\rho^{2}}+\sqrt{1+V_{\delta\gamma}(\delta,\gamma)}-2\,$, (59)
where
$V_{\delta\gamma}(\delta,\gamma)=\delta^{2}+k_{3}\left(\gamma+\frac{1}{2}\arctan(2k_{2}\delta)\right)^{2}\,$, (60)
with $k_{1},k_{2},k_{3}>0$ is a globally strict CLF on $\mathcal{S}$ for (2) with respect to the input pair $(v/\rho,\omega)$ in the sense of [39, Def. 2].
Example 4.4.
(Linear in $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$) Consider the quadratic cost function defined by
$\eta(r)=\frac{r^{2}}{2}\,$, (61)
with the derivative inverse $(\eta^{\prime})^{-1}(r)=r$ and the Legendre-Fenchel transform of $\ell\eta=\frac{r^{2}}{2}$. With this choice $\eta$, one finds the classical quadratic cost on both the state running cost and the control, as detailed in the basic quadratic inverse optimal controller (36) in Section 4.2, with control linear in $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$.
The simulation results for this controller with $\varepsilon_{1}=\varepsilon_{2}=1$, shown in Fig 4, has a clear contrast to the analytically derived controllers in [39] (see Fig. 4 in [39]). Notably, for many initial conditions, the controller chooses to reverse at first, adjusting its position to achieve a more favorable alignment before moving forward toward the target. This differs from the analytical controller, which tended to favor minimal deviation from the shortest path. Furthermore, for the initial condition where the unicycle starts parallel and directly above the target (cyan), the behavior resembles a conventional parallel parking maneuver (albeit reversed), more aligned with how a human driver would intuitively approach the task.
Figure 4: Trajectory of the inverse optimal controller designed using the CLF (59) with gains $(k_{1},k_{2},k_{3})=(6.5,3,7)$ and the quadratic control cost (61). The trajectory differences between Examples 4.44.7 are negligible and therefore omitted.
However, quadratic costs on control and state may be naive for nonlinear systems in general, particularly for the unicycle. Although they penalize input over the time horizon, they can still yield occasional large values outside actuator limits. As shown in Fig. 5(a), this results in forward velocity magnitudes up to about $10$ units/s and steering rate magnitudes up to about $2.5$ units/s—values not unreasonable but potentially beyond actuator range. In the next three examples, we present alternatives with stronger control penalties. Since the trajectories $y(x)$ remain nearly identical across cases, we focus instead on the differences in control effort.
(a) Control effort of the linearly proportional to $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$ controller (36).
(b) Control effort of the sublinearly proportional to $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$ controller (63).
Figure 5: Comparison in control effort between Example 4.4 and Example 4.5 with $\varepsilon_{1}=\varepsilon_{2}=1$.
Example 4.5.
(Sublinear/logarithmic in $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$) Consider the hyperbolic cosine cost function defined as
$\eta(r)=\cosh(r)-1\,$, (62)
with the associated derivative inverse $(\eta^{\prime})^{-1}(r)=\mathrm{arcsinh}(r)$ and the Legendre-Fenchel transformation $\ell\eta(r)=r\,\mathrm{arcsinh}(r)+1-\sqrt{1+r^{2}}$. The cost $\eta(r)=\cosh(r)-1$ imposes a larger penalty when the control effort is large as compared to the quadratic cost in Example 4.4. This is reflected in the minimizer feedback
$v^{*}$ $=-\rho\varepsilon_{1}\mathrm{arcsinh}\left(\varepsilon_{1}L_{\bar{g}_{1}}V\right)$ (63a)
$\omega^{*}$ $=-\varepsilon_{2}\mathrm{arcsinh}\left(\varepsilon_{2}L_{g_{2}}V\right)\,$, (63b)
where compared to the linear optimal controller in (36), the feedback law in (63) exhibits a sublinear dependence on $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$. Specifically, as $q=\varepsilon_{i}L_{g_{i}}V$ grows in magnitude, we observe that
${\mathrm{arcsinh}(q)}$ $={\ln\left(\sqrt{1+|q|^{2}}+|q|\right)}\mathrm{sgn}(q)$
$\approx{\ln(1+|q|)}\mathrm{sgn}(q)\,$, (64)
indicating that the feedback scales logarithmically with $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$. This implies that the controller grows more slowly than linearly with respect to $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$, and thus applies less aggressive control effort. The feedback in (63) can be interpreted as a less effort-intensive alternative to the linear controller in (36). This is reflected in Fig 5(b), with $\varepsilon_{1}=\varepsilon_{2}=1$, showing a nearly threefold reduction in control effort in the forward velocity input.
Example 4.6.
(Bounded/saturating control) While the hyperbolic cost on control effort significantly reduces the input magnitudes, one might still desire a stronger guarantee by directly enforcing a bound on the maximum allowable control input. To achieve bounded optimal control, we require that the inverse of $\eta^{\prime}$ in the general expression for the optimal control be a bounded function. For instance, setting $(\eta^{\prime})^{-1}(r)=\arctan(r)$ leads to
$\eta(r)=-\ln(\cos(r))\,$, (65)
which is of class $\mathcal{K}_{\infty}$ on the interval $0,\pi/2)$.
This choice implies that the control cost becomes unbounded as $|v|$ approaches the limit $\rho\varepsilon_{1}\pi/2$ and $|\omega|$ approaches the limit $\varepsilon_{2}\pi/2$, as seen in the optimal controller given by
$v^{*}$ $=-\rho\varepsilon_{1}\arctan\left(\varepsilon_{1}L_{\bar{g}_{1}}V\right)$ (66a)
$\omega^{*}$ $=-\varepsilon_{2}\arctan\left(\varepsilon_{2}L_{g_{2}}V\right)\,.$ (66b)
While the upper bound $|\omega|\leq\overline{\omega}$ is easily enforced by selecting $\varepsilon_{2}=2\overline{\omega}/\pi$, the upper bound $|v|\leq\overline{v}$ is more subtle. This is because the limiting value $\rho\varepsilon_{1}\pi/2$ depends on $\rho$. To address this, we define
$\varepsilon_{1}(\rho)=\frac{\overline{v}}{\sigma+\rho}\frac{2}{\pi}>0\,$, (67)
where $\sigma>0$ is a small constant, to ensure the bound holds. Fig 6(a) illustrates the result for $\bar{v}=\bar{\omega}=1$ and $\sigma=0.3$, showing that neither control inputs exceeds magnitude $1$.
The running cost on the state is then $\ell\eta(r)=r\arctan(r)-\frac{1}{2}\ln(1+r^{2})\approx\frac{\pi}{2}r$ as $r\rightarrow\infty$ indicating that the state incurs a linearly increasing penalty for large values of $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$, while remaining relatively tolerant near 0.
(a) Control effort of the bounded controller (66).
(b) Control effort of the bounded controller (69).
Figure 6: Comparison in control effort between Example 4.6 and Example 4.7 with $\bar{v}=\bar{\omega}=1$.
Example 4.7.
(“Relay-approximating” bounded control) Alternative to Example 4.6, consider the non-quadratic cost function defined by
$\eta(r)$ $=\frac{e}{e^{1/r}-e}\,$, (68)
which represents the control penalty, is infinitely flat near the origin and blows up at $r=1$. This means that small control magnitudes are penalized very lightly, while magnitudes approaching 1 incur an infinite cost, effectively enforcing a hard input constraint.
The state penalty $\ell\eta(r)$, while only numerically accessible, has infinite slope at $r=0$, indicating that any nonzero value of the state incurs a large cost. However, the function grows slowly for large arguments, revealing a cost structure that strongly discourages small errors but is relatively tolerant of large deviations.
The associated derivative inverse $(\eta^{\prime})^{-1}(r)=\frac{1}{1+\ln\left(1+\frac{1}{r}\right)}$ is continuous, monotonically increasing, bounded by 1, and has an infinite derivative at the origin. It closely resembles a smoothed approximation of the discontinuous signum function commonly used in sliding mode control. This behavior, characterized by high gain for small arguments and saturation for large ones, is typical of feedback laws arising in time-optimal control. Although the cost here is not explicitly time-optimal, it captures similar characteristics, including finite-time convergence under bounded input. Hence, the resulting optimal feedback law is given as
$v^{*}=-\rho\varepsilon_{1}\frac{\mathrm{sgn}(L_{\bar{g}_{1}}V)}{1+\ln\left(1+\dfrac{1}{\varepsilon_{1}|L_{\bar{g}_{1}}V|}\right)}$ (69a)
$\omega^{*}=-\varepsilon_{2}\frac{\mathrm{sgn}(L_{g_{2}}V)}{1+\ln\left(1+\dfrac{1}{\varepsilon_{2}|L_{g_{2}}V|}\right)}\,.$ (69b)
However, for similar reasons as in Example 4.6, to guarantee the upper bound $|v|\leq\overline{v}$ and $|\omega|\leq\overline{\omega}$, we choose
$\varepsilon_{1}(\rho)$ $=\frac{\overline{v}}{\sigma+\rho}>0\,$, (70)
for small $\sigma>0$ and $\varepsilon_{2}=\overline{\omega}$. Fig 6(b) shows the simulation with $\bar{v}=\bar{\omega}=1$ and $\sigma=0.3$. Notably, unlike the concentrated, high initial effort observed in Example 4.6, the control law (69) produces a more evenly sustained control effort over time. The settling times on the controls in Fig 6(a) and Fig 6(b) are different due to this more sustained control effort. This behavior is due to the cost function (68), which assigns minimal cost to small inputs, thereby tolerating low-magnitude control efforts over the trajectory. One can certainly not expect the achievement of fixed-time (FxT) stabilization under bounded control, but only finite-time (FT) stabilization.
Figure 7: Parallel parking trajectories for forward $L_{g}V$ control (red), reverse $L_{g}V$ control (blue), forward GloBa control [39, Thm. 7] (cyan), and reverse BoLSA control [39, Thm. 2] (yellow) with opacity representing different gain choices.
Gain margin and robustness to actuator saturation
Optimality is a benefit of inverse optimality. But the key practical benefit of the $L_{g}V$ inverse optimal control is the gain margin, i.e., the robustness to the uncertainty in the input coefficients. In the unicycle model, the coefficients on the inputs $(v,\omega)$ are taken, implicitly, as $(1,1)$. In the simplest of the inverse optimal controllers (36), the gains $\varepsilon_{1},\varepsilon_{2}>0$ are arbitrary, which means that the inverse optimal controller has a gain margin of $(0,\infty)$. This is a gain margin greater than the conventional gain margin $1/2,\infty)$ of optimal controllers for systems with drift, namely, the property that one can reduce the optimal controller’s gain by up to 2 and still preserve stability. The unusually strong gain margin $(0,\infty)$, which allows the reduction of the optimal controller’s gain for the unicycle to an arbitrarily low value is a consequence of the unicycle system being driftless.
Inverse optimality doesn’t afford only robustness to the uncertainty in the input coefficient. It also enables operation under arbitrarily low input saturation, as we observe in Examples 4.6 and 4.7. For example, taking $v^{*}$ as in (66), (67), we get the feedback
$v^{*}$ $=\bar{v}\ \underbrace{\frac{\rho}{\sigma+\rho}\frac{2}{\pi}\arctan\left(\frac{\bar{v}}{\sigma+\rho}\frac{2}{\pi}{\color[rgb]{0,0,1}L_{\bar{g}_{1}}V}\right)}_{\mbox{\small magnitude $<1$}}$ (71)
$\omega^{*}$ $=\bar{\omega}\ \underbrace{\frac{2}{\pi}\arctan\left(\bar{\omega}\frac{2}{\pi}{\color[rgb]{0,0,1}L_{g_{2}}V}\right)}_{\mbox{\small magnitude $<1$}}\,$, (72)
where $\bar{v},\bar{\omega}$ are the maximum forward and angular velocities, respectively, which the vehicle’s propulsion and steering actuators can deliver, which may be arbitrarily low. So, the feedback always acts in the right direction, which is determined by the sign of the expressions in blue under the $\arctan$. The $L_{\bar{g}_{1}}V$ and $L_{g_{2}}V$ terms in blue (with full expressions given in (36)) always deliver the “correct directions” for control by design: they are made to do so by our construction of the CLF $V(\rho,\delta,\gamma)$.
Nonholonomic parallel parking
Parallel parking has long been considered a benchmark maneuver for nonholonomic vehicles due to its intrinsic difficulty requiring precise coordination of orientation and position, often in confined spaces. It captures the core difficulty of planning and control for systems that cannot move freely in all directions, such as cars or mobile robots. With a wide range of control laws and infinitely many configurable parameters in both this paper and [39], our framework offers numerous options that can be tailored to the specific needs of a given task. Fig 7 highlights several representative examples from this toolbox. The $L_{g}V$ controllers exhibit behaviors that closely resemble human-like parallel parking, with the reverse $L_{g}V$ controller exactly representing the traditional approach used by drivers. In contrast, the standard GloBa and reversed BoLSA controllers [39, Thms. 2 and 7] produce smoother “S”-shaped trajectories that are more appropriate for maneuvering in tighter spaces.
5 Adaptive Stabilization under Unknown Input Coefficients
In view of the infinite gain margin as well as the $L_{g}V$ form feedback control law, in this section, we develop an adaptive control law for the unicycle model with model uncertainties.
Consider the model
$\dot{\rho}$ $=-b_{1}v\cos\gamma$ (73a)
$\dot{\delta}$ $=b_{1}\frac{v}{\rho}\sin(\gamma)$ (73b)
$\dot{\gamma}$ $=b_{1}\frac{v}{\rho}\sin(\gamma)-b_{2}\omega\,$, (73c)
where the constants $b_{1},b_{2}$ are positive but otherwise completely unknown. For example, $b_{1},b_{2}\in(0,1]$ may physically represent unknown wheel slippage coefficients (on a two-wheeled mobile robot).
Recall from (36) that
$L_{\bar{g}_{1}}V$ $:=-\dfrac{\partial V}{\partial\rho}\rho\cos\gamma+\left(\dfrac{\partial V}{\partial\delta}+\dfrac{\partial V}{\partial\gamma}\right)\sin\gamma$ (74a)
$L_{g_{2}}V$ $:=-\dfrac{\partial V}{\partial\gamma}\,$, (74b)
and introduce the adaptive control laws
$v$ $=$ $-\hat{\varepsilon}_{1}\rho L_{\bar{g}_{1}}V$ (75)
$\omega$ $=$ $-\hat{\varepsilon}_{2}L_{g_{2}}V\,$, (76)
where $\hat{\varepsilon}_{1},\hat{\varepsilon}_{2}$ are the online estimates of $1/b_{1},1/b_{2}$, respectively. Take any of our strict CLFs $V$, on any of the state spaces considered. The derivative of $V$ is
$\dot{V}=-(L_{\bar{g}_{1}}V)^{2}\left(1-b_{1}\tilde{\varepsilon}_{1}\right)-(L_{g_{2}}V)^{2}\left(1-b_{2}\tilde{\varepsilon}_{2}\right)\,$, (77)
where $\tilde{\varepsilon}_{i}=1/b_{i}-\hat{\varepsilon}_{i}$. Taking the adaptive CLF
$V_{\rm a}=\ln(1+n(V))+\frac{b_{1}}{2\mu_{1}}\tilde{\varepsilon}_{1}^{2}+\frac{b_{2}}{2\mu_{2}}\tilde{\varepsilon}_{2}^{2}\,$, (78)
with adaptation gains $\mu_{1},\mu_{2}>0$ and any normalization function $n\in\mathcal{K}_{\infty}\cap C^{1}$, including, for example, $n(V)=n_{0}V,n_{0}>0$, we get
$\dot{V}_{\rm a}=$ $-\frac{n^{\prime}(V)}{1+n(V)}\left[(L_{\bar{g}_{1}}V)^{2}+(L_{g_{2}}V)^{2}\right]$
$+b_{1}\tilde{\varepsilon}_{1}\left[\frac{n^{\prime}(V)}{1+n(V)}(L_{\bar{g}_{1}}V)^{2}-\frac{\dot{\hat{\varepsilon}}_{1}}{\mu_{1}}\right]$
$+b_{2}\tilde{\varepsilon}_{2}\left[\frac{n^{\prime}(V)}{1+n(V)}(L_{g_{2}}V)^{2}-\frac{\dot{\hat{\varepsilon}}_{2}}{\mu_{2}}\right]\,.$ (79)
Hence, we pick the update laws
$\dot{\hat{\varepsilon}}_{1}$ $=$ $\mu_{1}\frac{n^{\prime}(V)}{1+n(V)}(L_{\bar{g}_{1}}V)^{2}$ (80)
$\dot{\hat{\varepsilon}}_{2}$ $=$ $\mu_{2}\frac{n^{\prime}(V)}{1+n(V)}(L_{g_{2}}V)^{2}\,$, (81)
and obtain
$\dot{V}_{\rm a}=-\frac{n^{\prime}(V)}{1+n(V)}\left[(L_{\bar{g}_{1}}V)^{2}+(L_{g_{2}}V)^{2}\right]\,$, (82)
which is negative for all $(\rho,\delta,\gamma)\neq(0,0,0)$ on the state space considered but is not negative definite in the overall state of the adaptive system, which also includes $(\tilde{\varepsilon}_{1},\tilde{\varepsilon}_{2})$. With LaSalle’s theorem, we obtain the following result.
Theorem 5.8.
Consider the system (73) with arbitrary unknown $b_{1},b_{2}>0$, along with the control law (75), (76) and the update laws (80), (81). For each CLF $V$ given by [39, Thms. 1-8], with class $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2}$ such that
$\alpha_{1}\left(|\rho,\delta,\gamma|_{\mathcal{Q}}\right)\leq V(\rho,\delta,\gamma)\leq\alpha_{2}\left(|\rho,\delta,\gamma|_{\mathcal{Q}}\right)\,$, (83)
for all initial conditions $(\rho(0),\delta(0),\gamma(0))$ on its respective state space $\mathcal{Q}\in\{\mathcal{S},\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3}\}$, and for all parameter initial conditions $\hat{\varepsilon}_{1}(0)\in\mathbb{R},\hat{\varepsilon}_{2}(0)\in\mathbb{R}$, the following holds:
$\Upsilon(t)\leq a_{1}^{-1}\left(M\left(e^{m\,a_{2}(\Upsilon(0))}-1\right)\right)\,,\quad\forall t\geq 0\,$, (84)
with
$\Upsilon$ $=|(\rho,\delta,\gamma)|_{\mathcal{Q}}+|(\tilde{\varepsilon}_{1},\tilde{\varepsilon}_{2})|\,$, (85)
where
$a_{1}(r)$ $=$ $\min\{n\circ\alpha_{1}(r),r^{2}\}$ (86)
$a_{2}(r)$ $=$ $\max\{n\circ\alpha_{2}(r),r^{2}\}$ (87)
are $\mathcal{K}_{\infty}$ functions and
$M=\max\left\{1,\dfrac{c_{2}}{c_{1}}\right\}\,$, $\quad m=\max\left\{c_{2},\frac{1}{c_{2}}\right\}\,$, (88)
$c_{1}=\min\left\{\frac{b_{1}}{2\mu_{1}},\frac{b_{2}}{2\mu_{2}}\right\}\,$, $\quad c_{2}=\max\left\{\frac{b_{1}}{2\mu_{1}},\frac{b_{2}}{2\mu_{2}}\right\}\,.$ (89)
In addition, $\rho(t),\delta(t),\gamma(t)\rightarrow 0$ as $t\rightarrow\infty$.
Proof 5.9.
For all $t\geq 0$, from (82) it follows that $V_{a}(t)\leq V_{a}(0)$, and from (78) that
$\ln\bigl(1+n\left(V(t)\right)\bigr)+c_{1}\left(\tilde{\varepsilon}_{1}^{2}(t)+\tilde{\varepsilon}_{2}^{2}(t)\right)$
$\leq\ln\bigl(1+n\left(V(0)\right)\bigr)+c_{2}\left(\tilde{\varepsilon}_{1}^{2}(0)+\tilde{\varepsilon}_{2}^{2}(0)\right)\,.$ (90)
Further calculations lead to (84) with (88), (89). The convergence follows from (82) with LaSalle’s theorem.
(a) Norm of $\Xi\coloneqq(\rho,\delta,\gamma)$ over time for each control law showing faster decay rate for both adaptive control laws.
(b) Maximum magnitude of the control input is smaller for both adaptive control laws compared to the $L_{g}V$ controller despite the faster decay rate.
(c) $\hat{\varepsilon}_{1}$ and $\hat{\varepsilon}_{2}$ grows above 11 leading to the faster decay rate.
Figure 8: Comparison of state trajectories and control inputs under the quadratic cost $L_{g}V$ controller with both $b_{1}\varepsilon_{1}$ and $b_{2}\varepsilon_{2}$ taken as 11 and the proposed adaptive controller with two different adaptation gains.
To illustrate the performance of the proposed method, we consider the BoFo CLF $V=\rho^{2}+V_{\delta\gamma}$, where $V_{\delta\gamma}$ is defined in [39, Thm. 6], and choose gains $k_{1}=3$, $k_{2}=3.5$, and $k_{3}=4$. In Fig. 8, we compare the quadratic cost $L_{g}V$ controller (red), given by $v=-\rho L_{\bar{g}_{1}}V$ and $\omega=-L_{g_{2}}V$, with the adaptive control law in (80)–(81), which uses a normalization function of the form $n(V)=V$, initial parameter estimates $\hat{\varepsilon}_{1}(0)=\hat{\varepsilon}_{2}(0)=0$, and adaptation gains $\mu_{1}=\mu_{2}=0.5$ (blue) and $\mu_{1}=\mu_{2}=1$ (cyan). The system is initialized at $(\rho(0),\delta(0),\gamma(0))=(1,-\pi/2,-\pi/2)$, under uncertainty in $b_{1}=b_{2}=1$. For the quadratic cost $L_{g}V$ controller, this corresponds to the assumption of $b_{1}\varepsilon_{1}=b_{2}\varepsilon_{2}=1$. The adaptive controller exhibited smaller peak values in both forward velocity and steering input (Figure 8(b)), even while achieving a faster decay of the state norm (Figure 8(a)). This behavior is clearly due to $\hat{\varepsilon}_{1,2}$ starting small but quickly growing beyond $1$, thereby accelerating the decay rate to be faster than the $L_{g}V$ controller.
The adaptive control law (75), (76), (80), (81) has similarities with the inverse optimal controller (36). The difference is that the adaptive controller works even if it starts with gains of the wrong sign, $\hat{\varepsilon}_{1}(0)<0,\hat{\varepsilon}_{2}(0)<0$. Additionally, the adaptive controller’s gain is improved, online, by learning from the transients reflected in $L_{\bar{g}}V$.
Of course, all update laws exhibit drift in the presence of disturbances. By adding leakage to the update laws (80) and (81), as is standard, boundedness of the vehicle states and parameter estimates, as well as practical regulation of $\rho(t),\delta(t),\gamma(t)$, would be ensured.
6 Prescribed and Fixed-Time Nonholonomic Parking
In this section, we modify the controllers designed in Section 3 such that the states $(\rho,\delta,\gamma)$ converge to the point $\rho=\delta=\gamma=0$ in user-defined prescribed time $T>0$.
(a) Control inputs
(b) State trajectories
Figure 9: PT-parking and FxT-parking using the control laws (91) and (101) for $T=2$ ($p=0.3$ in the FxT case). The control gains are $k_{1},k_{2},k_{3},k_{4}]=[1,2.2,2.5,0.5]$ for the PT case and $k_{1},k_{2},k_{3},k_{4}]=[3,0.8,0.5,1]$ for the FxT case, applied to both the unidirectional BAR-FLi controller (7)–(9) and the bidirectional backstepping controller (21)–(23). Different gains are chosen to achieve comparable behaviors given the distinct scaling factors inherent in the PT and FxT formulations.
Theorem 6.10 (PT-Parking).
Consider the system (2) in closed-loop with
$v(t)$ $=(1+\nu^{2}(t-t_{0}))\tilde{v}(t)$ (91a)
$\omega(t)$ $=(1+\nu^{2}(t-t_{0}))\left(\frac{v}{\rho}\sin(\gamma)+\tilde{\omega}(t)\right)$ (91b)
where
$\nu(t-t_{0})=\tan\left(\frac{\pi(t-t_{0})}{2T}\right)$ (92)
with $T>0$ and the forward and angular velocities $\tilde{v},\tilde{\omega}$ are respectively chosen from: (i) (7) and (9); or (ii) (21) and (23), each of which requires the gains to satisfy $k_{1},k_{2},k_{3},k_{4}>0$. Then,
  1. 1.
    there exists $K,\lambda>0$, such that $|(\rho(t),\delta(t),\gamma(t))|_{\mathcal{Q}}\leq K|(\rho_{0},\delta_{0},\gamma_{0})|_{\mathcal{Q}}{\rm e}^{-\frac{2T}{\pi}\lambda\nu(t-t_{0})}$ for all $t\in[t_{0},t_{0}+T)$, where the state-space $\mathcal{Q}$ corresponds to any of $\mathcal{S},\mathcal{S}_{2}$ depending on the choice of $\tilde{v}$, $\tilde{\omega}$,
  2. 2.
    and the input signals in the control laws (91) are bounded and converge to zero as $t\rightarrow t_{0}+T$.
Proof 6.11.
Step 1: We dilate the finite time interval $t\in[t_{0},t_{0}+T)$ to the infinite time interval $\tau\in[t_{0},\infty)$ via the transformation $t\rightarrow\tau$:
$\tau=t_{0}+\frac{2T}{\pi}\nu(t-t_{0}).$ (93)
Considering the fact that
$\frac{{\rm d}t}{{\rm d}\tau}=\frac{1}{1+\left(\frac{\pi}{2T}(\tau-t_{0})\right)^{2}}\,$, (94)
the control laws (91) in time $\tau$ are
$v(\tau)$ $=\left(1+\frac{\pi^{2}}{4T^{2}}(\tau-t_{0})^{2}\right)\tilde{v}$, (95a)
[…]
Hence, the closed-loops in time $\tau$ are exactly the same as (16), (17), (18) with (7) and (9) or as (28) (29), (30) with (21) and (23) where ${\rm d}/{\rm d}t$ replaced by ${\rm d}/{\rm d}\tau$. Thus, the corresponding stability guarantees stated in Theorems 12 hold in time $\tau$, respectively. This implies the existence of positive constants $K,\lambda>0$, such that,
$\hskip-10.81218pt\left\lvert(\rho(\tau),\delta(\tau),\gamma(\tau)\right\rvert_{\mathcal{Q}}\leq K\left\lvert(\rho(t_{0}),\delta(t_{0}),\gamma(t_{0})\right\rvert_{\mathcal{Q}}{\rm e}^{-\lambda(\tau-t_{0})}$ (96)
where $\mathcal{Q}$ corresponds to any of $\mathcal{S},\mathcal{S}_{2}$ depending on the choice of $\tilde{v}$ and $\tilde{\omega}$. In addition, we bound the control laws (95) as
$\left\lvert v(\tau)\right\rvert+\left\lvert\omega(\tau)\right\rvert\leq$ $\left(1+\frac{\pi^{2}}{4T^{2}}(\tau-t_{0})^{2}\right)\times$ (97)
$\tilde{\alpha}(\left\lvert(\rho(\tau),\delta(\tau),\gamma(\tau))\right\rvert_{\mathcal{Q}}),$
where $\tilde{\alpha}\in\mathcal{K}_{\infty}$ is defined according to the chosen controller and the corresponding state space $\mathcal{Q}$ as follows: On $\mathcal{Q}=\mathcal{S}$ the function $\tilde{\alpha}(r)=\max\{c_{1},c_{2}\}(r+r^{2}+r^{3})$ corresponds to (21), (23) with $c_{1}=2k_{1}\sqrt{1+k_{2}^{2}}+k_{4}(1+k_{2})+k_{3}\sqrt{1+k_{2}^{2}}+k_{1}k_{2}(\sqrt{1+k_{2}^{2}}(1+k_{2})+1)$ and $c_{2}=4k_{1}k_{2}+k_{3}+4k_{2}k_{3}+2k_{1}k_{2}^{2}(1+k_{2})$, while on $\mathcal{Q}=\mathcal{S}_{2}$, the function $\tilde{\alpha}(r)=c_{1}(r+r^{2}+r^{3})$ corresponds to (7), (9). Based on (96) and (97), we obtain
$\left\lvert v(\tau)\right\rvert$+ $\left\lvert\omega(\tau)\right\rvert\leq\left(1+\frac{\pi^{2}}{4T^{2}}(\tau-t_{0})^{2}\right)\times$ (98)
$\tilde{\alpha}\left(K\left\lvert\left(\rho(t_{0}),\delta(t_{0}),\gamma(t_{0})\right)\right\rvert_{\mathcal{Q}}{\rm e}^{-\lambda(\tau-t_{0})}\right).$
Considering the functions $\tilde{\alpha}$ as in (97), established above to all satisfy $\tilde{\alpha}(r)\leq a(r+r^{2}+r^{3}),a\in\mathbb{R}_{>0}$, and letting $\tau\rightarrow\infty$ in (98), we observe that $\left\lvert v(\tau)\right\rvert+\left\lvert\omega(\tau)\right\rvert$ converges to zero.
Step 2: We now convert the above results into time $t$ by contracting the infinite time interval $t_{0},\infty)$ into the finite time interval $t_{0},t_{0}+T)$ in time $t$, by taking the inverse transformation of (93) as
$t=t_{0}+\frac{2T}{\pi}\arctan\left(\frac{\pi}{2T}(\tau-t_{0})\right).$ (99)
Then, (96) in time $t$ reads
$\left\lvert(\rho(\tau),\delta(\tau),\gamma(\tau)\right\rvert_{\mathcal{Q}}\leq K\left\lvert(\rho(t_{0}),\delta(t_{0}),\gamma(t_{0})\right\rvert_{\mathcal{Q}}{\rm e}^{-\frac{2T}{\pi}\lambda\nu(t-t_{0})}$ (100)
which proves property 1 of Theorem 6.10. Analogously, (98) in time $t$ reads as
$\left\lvert v(t)\right\rvert+\left\lvert\omega(t)\right\rvert\leq(1+\nu^{2}(t-t_{0}))\times$
$\tilde{\alpha}\left(K\left\lvert(\rho(t_{0}),\delta(t_{0}),\gamma(t_{0})\right\rvert_{\mathcal{Q}}{\rm e}^{-\frac{2T}{\pi}\lambda\nu(t-t_{0})}\right).$
Letting $t\rightarrow t_{0}+T$ in (6.11), we conclude that $\left\lvert v(t)\right\rvert+\left\lvert\omega(t)\right\rvert$ is bounded and converges to zero. Thus, property 2 of Theorem 6.10 holds.
Theorem 6.12 (FxT-Parking).
Consider the system (2) in closed-loop with
$v(t)$ $=\kappa(\rho,\delta,\gamma)\tilde{v}(t)$ (101a)
$\omega(t)$ $=\kappa(\rho,\delta,\gamma)\left(\frac{v}{\rho}\sin(\gamma)+\tilde{\omega}(t)\right)$ (101b)
where
$\kappa(\rho,\delta,\gamma)=\frac{1}{\underline{c}pT}\exp(V^{p})V^{-p},$ (102)
with $p\in(0,\frac{1}{2})$ and $T>0$ and the forward and angular velocities $\tilde{v},\tilde{\omega}$, the backstepping transformations $z$ and the Lyapunov functions $V$ are respectively chosen from: (i) (7), (9), (13) and (15) with $\underline{c}$ as in (20); or (ii) (21), (23), (25) and (27) with $\underline{c}$ as in (32), each of which requires the gains to satisfy $k_{1},k_{2},k_{3},k_{4}>0$. Then,
  1. 1.
    there exists $K_{1},K_{2}>0$, such that it holds
    $\hskip-11.38092pt\begin{aligned} |(&\rho(t),\delta(t),\gamma(t))|_{\mathcal{Q}}\leq\\ &K_{1}\left[\ln\left(\frac{1}{\frac{t-t_{0}}{T}+\exp(-K_{2}\left\lvert(\rho_{0},\delta_{0},\gamma_{0})\right\rvert^{2p}_{\mathcal{Q}})}\right)\right]^{\frac{1}{2p}}\end{aligned}$ (103)
    for all $t\in[t_{0},t_{0}+T_{\rm s}(\rho_{0},\delta_{0},\gamma_{0}))$, and $|(\rho(t),\delta(t),\gamma(t))|_{\mathcal{Q}}=0$ for all $t\geq t_{0}+T_{\rm s}(\rho_{0},\delta_{0},\gamma_{0})$, and for all $(\rho_{0},\delta_{0},\gamma_{0})$, the settling time $T_{\rm s}$ is such that
    $\hskip-5.69046ptT_{\rm s}(\rho_{0},\delta_{0},\gamma_{0})\leq[1-\exp(-K_{2}\left\lvert(\rho_{0},\delta_{0},\gamma_{0})\right\rvert_{\mathcal{Q}}^{2p})]T\leq T\,$, (104)
    where the state-space $\mathcal{Q}$ corresponds to either $\mathcal{S}$ or $\mathcal{S}_{2}$ depending on the choice of $\tilde{v}$ and $\tilde{\omega}$,
  2. 2.
    and the input signals in the control laws (101) are bounded and converge to zero no later than $t=t_{0}+T$.
Proof 6.13.
With the rescaled control laws (101), we follow the proofs of Theorems 1 and 2 and arrive at the rescaled versions of (20) and (32), which now read as
$\dot{V}\leq-\underline{c}k(\rho,\delta,\gamma)V=-\frac{1}{pT}\exp(V^{p})V^{1-p}.$ (105)
By the comparison principle, this implies that
$V(t)\leq\left[\ln\left(\frac{1}{\frac{t-t_{0}}{T_{c}}+\exp(-V_{0}^{p})}\right)\right]^{1/p}\,.$ (106)
Considering the fact that both Lyapunov functions (15) and (27) are bounded as
$k_{1}\left\lvert(\rho,\delta,\gamma)\right\rvert_{\mathcal{Q}}^{2}\leq V\leq k_{2}\left\lvert(\rho,\delta,\gamma)\right\rvert_{\mathcal{Q}}^{2}\,$, (107)
where $k_{1}=\min\left\{1/2,k_{1}/(2k_{1}k_{2}^{2}+k_{3})\right\}$, $k_{2}=\max\left\{1,1+2k_{1}k_{2}^{2}/k_{3},2k_{1}/k_{3}\right\}$, on their respective spate-space $\mathcal{Q}=\{\mathcal{S},\mathcal{S}_{2}\}$, we arrive at (103). The norm of the control laws (101) is
$\left\lvert v\right\rvert+\left\lvert\omega\right\rvert\leq\frac{1}{\underline{c}pT}\exp(V^{p})V^{-p}\tilde{\alpha}(\left\lvert(\rho,\delta,\gamma\right\rvert_{\mathcal{Q}})\,$, (108)
where $\tilde{\alpha}(r)=a(r+r^{2}+r^{3}),a\in\mathbb{R}_{>0}$ as already established in (97). Considering the lower bound from (107), we obtain
$\left\lvert v\right\rvert+\left\lvert\omega\right\rvert\leq$ $\frac{a\exp(V^{p})}{\underline{c}k_{1}^{p}pT}(\left\lvert(\rho,\delta,\gamma)\right\rvert_{\mathcal{Q}}^{1-2p}$ (109)
$+\left\lvert(\rho,\delta,\gamma)\right\rvert_{\mathcal{Q}}^{2-2p}+\left\lvert(\rho,\delta,\gamma)\right\rvert_{\mathcal{Q}}^{3-2p})$
where 12p>01-2p>0, 22p>02-2p>0, 32p>03-2p>0 since $p\in(0,1/2)$, hence (109) is bounded as $\left\lvert(\rho,\delta,\gamma)\right\rvert_{\mathcal{Q}}$ goes to zero. Now, consider (103) and note that if $(t-t_{0})/T+\exp(-K_{2}\left\lvert(\rho,\delta,\gamma)\right\rvert_{\mathcal{Q}}^{2p})=1$, then $\left\lvert(\rho,\delta,\gamma)\right\rvert_{\mathcal{Q}}=0$. Since $\exp$” is less than or equal to 1, the settling time $T_{s}(\rho_{0},\delta_{0},\gamma_{0})\leq T[1-\exp(-K_{2}\left\lvert(\rho_{0},\delta_{0},\gamma_{0})\right\rvert_{\mathcal{Q}}^{2p})]\leq T$, from which statements 1 and 2 of Theorem 6.12 follow.
The PT-controllers (91) admit a directly tunable settling time, since the convergence horizon $T$ can be explicitly prescribed by the designer and remains independent of both the initial conditions and the controller gains. In contrast, the more conservative FxT-controllers (101) provide only an upper-bound estimate of the settling time of $(\rho,\delta,\gamma)$ to zero given in (104) that depends on the initial conditions and the chosen control parameters. Consequently, the designer can specify only an overestimate $T$ of the true convergence time. Moreover, the FxT-control laws (101) are nonsmooth even in the polar coordinates and can add additional discontinuities to the control laws when transformed to Cartesian space, making their implementation more sensitive to measurement noise. As opposed to that, the PT-control laws (91) yield control signals that are smooth in both time and the states $(\rho,\delta,\gamma)$.
The PT-controller is meaningful only for the time interval $t_{0},t_{0}+T)$, however, this feature could be advantageous in applications that require the motion to be completed within a prescribed-time window, after which the control can be safely deactivated. In contrast, the FxT controller operates continuously beyond the convergence to zero, which makes it more suitable for ongoing parking tasks where an infinite horizon is necessary. In both cases, the control inputs remain bounded and converge smoothly to zero, ensuring comparable terminal behavior despite the structural differences between the two approaches. The properties asserted by Theorems 6.10 and 6.12 are illustrated by the simulations shown in Fig. 9.
7 Nonovershooting Control: Strictly Forward Parking without Curb Violation
Finally, we incorporate an analytical nonovershooting formulation that enforces safety constraints. Suppose the goal is to regulate the unicycle to the point $x=y=\theta=0$, but without overshooting in $y$, i.e., with $y(t)\leq 0$ for all finite time—starting and remaining in the lower half of the Cartesian plane, akin to parking a vehicle without going over the curb. The full interval in which $y=-\rho\sin\delta\leq 0$ where $\delta\in[-\pi,\pi)$ corresponds to $\delta\in[0,\pi)$. So, we formulate our “nonovershooting control” as the stabilization of $\rho=\delta=\gamma=0$ with an arbitrary $\delta_{0}\in(0,\pi)$ and with $\gamma_{0}$ possibly limited to some interval around $\gamma=0$.
We utilize the BAR-FLi CLFs as in [39, Thm. 8], which have barrier-like properties as discussed in [39, Section III.D] and serve to restrict the polar angle $\delta$. Building on the backstepping transformation of BAR-FLi, we design $\tilde{\omega}(t)$ to ensure that $\delta(t)\in[0,\pi)$ for all $t\geq 0$, which implies that $y(t)\leq 0$ for all $t\geq 0$. The following result introduces a control law that renders the origin GAS on $\mathcal{S}_{2}$, and further ensures $\delta(t)\in[0,\pi)$ for all $t\geq 0$ if the initial condition satisfies $\delta_{0}\in(0,\pi)$ and $\gamma_{0}\in(-\pi/4,\pi/4)$ through an appropriate gain selection. Furthermore, under the same initial condition constraints, the forward velocity input $v(t)$ remains nonnegative for all $t\geq 0$, which is crucial in applications where the unicycle model represents fixed-wing aircraft or guided missiles. Fig 10 illustrates the corresponding admissible region in the $(\delta,\gamma)$ space and its geometric interpretation in Cartesian coordinates.
Theorem 7.14.
For the system (2), the control laws
$v$ $=$ $k_{1}\rho\cos(\gamma)$ (110)
$\omega$ $=$ $\dfrac{k_{1}}{2}\sin(2\gamma)+\tilde{\omega}$ (111)
$\tilde{\omega}$ $=$ $\left(k_{4}+\dfrac{k_{3}}{k_{2}}\psi^{2}(z,\delta)N(\delta)\left(1+\tan^{2}\displaystyle\frac{\delta}{2}\right)\right)z$ (112)
$+\dfrac{k_{1}}{2}\dfrac{k_{2}}{\cos^{2}\dfrac{\delta}{2}+16k_{2}^{2}\sin^{2}\dfrac{\delta}{2}}\sin(2\gamma)\,$,
where
$z$ $=$ $\gamma+\frac{1}{2}\arctan\left(4k_{2}\tan\frac{\delta}{2}\right)$ (113)
$N(\delta)$ $=$ $\sqrt{1+16k_{2}^{2}\tan^{2}(\delta/2)}$ (114)
$\psi(z,\gamma)$ $=$ $\frac{\sin(2z-2\gamma)+\sin(2\gamma)}{2z}\,$, (115)
and with $k_{1},k_{2},k_{3},k_{4}>0$, render the point $\rho=\delta=\gamma=0$ GAS on $\mathcal{S}_{2}$, in accordance with [39, Def. 1]. Moreover, if the initial conditions satisfy $\delta_{0}\in(0,\pi)$ and $\gamma_{0}\in(-\pi/4,\pi/4)$ with $k_{2}$ chosen as
$k_{2}\in\left(\frac{\tan\left(\max\{0,-2\gamma_{0}\}\right)}{4\tan(\delta_{0}/2)},\frac{\tan\left(\frac{\pi}{2}+\min\{0,-2\gamma_{0}\}\right)}{4\tan(\delta_{0}/2)}\right)\,$, (116)
the polar angle $\delta(t)$ remains in the interval $0,\pi)$, implying that $(x(t),y(t),\theta(t))\rightarrow 0$ as $t\rightarrow\infty$ with $y(t)\leq 0$, and the forward velocity input $v(t)$ remains nonnegative for all $t\geq 0$.
Proof 7.15.
Consider again the BAR-FLi Lyapunov function introduced in [39, Thm. 8],
$V_{\delta\gamma}(\delta,\gamma)=4\tan^{2}\frac{\delta}{2}+q^{2}z^{2}\,$, (117)
where $z$ is defined in (113) and $q=\sqrt{k_{1}/k_{3}}$. The time derivative of (117) with (110) and (111) yields
$\dot{V}_{\delta\gamma}=$ $-\frac{8k_{1}k_{2}\tan^{2}(\delta/2)}{N(\delta)}\left(1+\tan^{2}\displaystyle\frac{\delta}{2}\right)$
$+2q^{2}\,z\biggl[\frac{k_{1}k_{2}\sin(2\gamma)}{2\cos^{2}(\delta/2)+32k_{2}^{2}\sin^{2}(\delta/2)}$
$+2k_{3}\psi(z,\gamma)\left(1+\tan^{2}\frac{\delta}{2}\right)\tan\frac{\delta}{2}-\tilde{\omega}\biggr$]. (118)
Noting that $N(\delta)>0$, from Young’s inequality, we get
$4k_{1}z\psi(z,\gamma)\tan(\delta/2)$
$\qquad\leq\frac{2k_{1}k_{2}\tan^{2}(\delta/2)}{N(\delta)}+\frac{2k_{1}}{k_{2}}N(\delta)\psi^{2}(z,\delta)z^{2}\,.$ (119)
Then,
$\dot{V}_{\delta\gamma}\leq$ $-\frac{6k_{1}k_{2}\tan^{2}(\delta/2)}{N(\delta)}\left(1+\tan^{2}\displaystyle\frac{\delta}{2}\right)$
$+\frac{2k_{1}}{k_{2}}N(\delta)\psi^{2}(z,\delta)\left(1+\tan^{2}\displaystyle\frac{\delta}{2}\right)z^{2}$
$+2q^{2}z\Biggl[\frac{k_{1}k_{2}\sin(2\gamma)}{2\cos^{2}(\delta/2)+32k_{2}^{2}\sin^{2}(\delta/2)\bigr)}-\tilde{\omega}\Biggr$]\,. (120)
Choosing the $\tilde{\omega}$ as (112) we get
$\dot{V}_{\delta\gamma}\leq-\frac{6k_{1}k_{2}}{N(\delta)}\left(1+\tan^{2}\displaystyle\frac{\delta}{2}\right)\tan^{2}\frac{\delta}{2}-2k_{4}q^{2}z^{2}\,.$ (121)
Thus, under the control laws (110) and (111) with (112), the time derivative of the CLF
$V(\rho,\delta,\gamma)=\rho^{2}+4\tan^{2}\frac{\delta}{2}+q^{2}z^{2}\,$, (122)
is negative definite on $\{\rho\geq 0\}\times\mathcal{T}_{2}$. By reasoning similar to that of the proof of [39, Thm. 8], the point $\rho=\delta=\gamma=0$ is GAS on $\mathcal{S}_{2}$, implying that $(x(t),y(t),\theta(t))\to 0$ as $t\to\infty$.
To prove the second part, first observe that with (111) and (112) the time derivative of the backstepping transformation (113) yields
$\dot{z}=-\left(k_{4}+\dfrac{k_{3}}{k_{2}}{\psi^{2}(z,\delta)}N(\delta)\left(1+\tan^{2}\displaystyle\frac{\delta}{2}\right)\right)z\,$, (123)
which guarantees $z(t)\geq 0$ and is decreasing for all $z_{0}\geq 0$. Then, if $\delta_{0}\in(0,\pi)$, $\gamma_{0}\in(-\pi/4,\pi/4)$, and $k_{2}$ is chosen as in (116), the expression (113) guarantees that $z_{0}\in(0,\pi/4)$ and hence $z(t)\in[0,\pi/4)$ for all $t\geq 0$. Given that, therefore, $\frac{k_{1}\cos(2z(t))}{2N(\delta(t))}\tan(2z(t))\geq 0$ and rewriting $\dot{\delta}$ as
$\dot{\delta}$ $=\frac{k_{1}\cos(2z)}{2N(\delta)}\left(\tan(2z)-4k_{2}\tan\frac{\delta}{2}\right)\,$, (124)
we observe that $\dot{\delta}\geq-\frac{2k_{1}k_{2}}{N(\delta)}\tan(\delta/2)$ which for $\delta=0$ implies $\dot{\delta}\geq 0$. Hence, $\delta(t)$ remains nonnegative for all $t\geq 0$. Additionally, from (124), we observe that whenever $\delta(t)\geq 2\arctan\left(\frac{\tan(2z(t))}{4k_{2}}\right)$ for any $t\geq 0$, it follows that $\dot{\delta}(t)\leq 0$ and since $z(t)<\pi/4$ for all $t\geq 0$, $\delta(t)<2\arctan(\infty)=\pi$ holds and ensures that $\delta(t)<\pi$ for all $t\geq 0$ if $\delta_{0}\in(0,\pi)$. Thus, $\delta(t)\in[0,\pi)$ for all $t\geq 0$, which implies that $y(t)\leq 0$ for all $t\geq 0$. Finally, since $z(t)\in[0,\pi/4)$ and $\delta(t)\in[0,\pi)$ for all $t\geq 0$, then by (113) it follows that $\gamma(t)\in(-\pi/4,\pi/4)$, yielding $\cos(\gamma(t))>\sqrt{2}/2$ and thus guarantees that the forward velocity control law (110) remains nonnegative for all $t\geq 0$.
(a) Admissible set in $(\gamma,\delta)$
(b) Geometry of the LOS angle $\gamma$
Figure 10: Visualization of the angular variables used in the nonovershooting controller design. (a) The admissible set in $\delta$ and $\gamma$ for guaranteed nonovershoot. (b) Illustration of the admissible LOS angle $\gamma$ for selected initial positions relative to the target’s position and heading (blue).
To demonstrate this result, Fig 11 presents simulations that illustrate its ability to enforce the curb constraint under a range of initial conditions. The strict enforcement of the safety constraint is verified in Fig 12(a), where $\delta(t)\geq 0$ holds for all initial conditions and time, ensuring the system remains within the lower half-plane. Notably, in the extreme case where the initial condition satisfies $\delta_{0}\approx\pi$ (grey trajectory), the controller actively suppresses the growth of $\delta(t)$ to prevent any potential violation of the safety constraint.
Figure 11: Trajectories under the nonovershooting controls (110) and (111), using (112) with gains $k_{1}=k_{3}=k_{4}=1$, and $k_{2}$ chosen within the interval in (116).
(a) $\delta(t)$ remains non-negative for all time, with $\gamma(t)$ exhibiting abrupt turns in certain trajectories (blue, cyan, and lavender) once $\delta(t)$ reaches zero.
(b) Control inputs $v$ and $\omega$ for the nonovershooting trajectories.
Figure 12: Simulation results under nonovershooting control laws.
For several initial conditions (blue, cyan, and lavender) one observes an abrupt turn in the trajectory once the unicycle reaches the boundary, represented as a sharp jump in $\gamma(t)$ in Fig 12(a), akin to the discontinuity typical of conventional safety filters. The corresponding impulse-like spikes in $\omega(t)$ (Fig 12(b)) are not due to controller discontinuity, but a result of the necessary design choice of large $k_{2}$ values for these cases ($392.85$, $221.09$, and $127.65$, respectively). Recalling the expression for $\tilde{\omega}$ in (112), when $k_{2}$ is large, the last term remains small for large $\delta$ since the $k_{2}^{2}$ term in the denominator dominates the $k_{2}$ term in the numerator. As $\delta$ approaches zero, the denominator becomes small, causing the expression to grow rapidly. However, once the unicycle reaches the boundary and $\gamma(t)$ becomes zero, the term vanishes. This results in a control input that appears impulse-like and, while it can be large, it is still continuous.
8 Conclusion
This paper presents constructive analytical results for nonmodular exponential stabilization, inverse optimality, robustness, adaptive control, prescribed/fixed-time convergence, and safety of the unicycle system. We develop nonmodular backstepping feedback laws that guarantee global exponential stability with either unidirectional or bidirectional velocity actuation. Leveraging families of strict CLFs, we introduce a general inverse optimal framework that yields broad families of stabilizers with infinite gain margins, ensuring robustness under arbitrarily low input saturation and actuator uncertainty. To further address uncertainty, we design adaptive laws for unknown input coefficients that achieve improved transient performance with reduced initial control effort. We further present controllers that ensure stabilization within a user-defined time horizon through two systematic approaches: a time-dilated prescribed-time design that guarantees smooth convergence of both states and inputs to zero, and a homogeneity-based fixed-time control scheme that offers an analytical bound on the settling time. Finally, nonovershooting control ensures geometric safety constraints and strictly forward motion without optimization-based enforcement. Together with Part I, this work establishes a broad constructive foundation for the stabilization of the nonholonomic unicycle, unifying multiple advanced control properties in a systematic and analytically transparent framework.